# I/O in Machine Learning Applications on HPC Systems: A 360-degree Survey

### v1

@@ Acknowledgements.
- Eisenman et al. (2022) Assaf Eisenman, Kiran Kumar Matam, Steven Ingram, Dheevatsa Mudigere, Raghuraman Krishnamoorthi, Krishnakumar Nair, Misha Smelyanskiy, and Murali Annavaram. 2022. Check-N-Run: a Checkpointing System for Training Deep Learning Recommendation Models. In 19th USENIX Symposium on Networked Systems Design and Implementation (NSDI 22). USENIX Association, Renton, WA, 929–943. https://www.usenix.org/conference/nsdi22/presentation/eisenman
@@
- Hu et al. (2021) Weihua Hu, Matthias Fey, Marinka Zitnik, Yuxiao Dong, Hongyu Ren, Bowen Liu, Michele Catasta, Jure Leskovec, Regina Barzilay, Peter Battaglia, Yoshua Bengio, Michael Bronstein, Stephan Günnemann, Will Hamilton, Tommi Jaakkola, Stefanie Jegelka, Maximilian Nickel, Chris Re, Le Song, Jian Tang, Max Welling, and Rich Zemel. 2021. Open Graph Benchmark: Datasets for Machine Learning on Graphs Steering Committee. https://ogb.stanford.edu.
- Jain et al. (2020) Abhinav Jain, Hima Patel, Lokesh Nagalapatti, Nitin Gupta, Sameep Mehta, Shanmukha Guttula, Shashank Mujumdar, Shazia Afzal, Ruhi Sharma Mittal, and Vitobha Munigala. 2020. Overview and Importance of Data Quality for Machine Learning Tasks. In Proceedings of the 26th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining (Virtual Event, CA, USA) (KDD ’20). Association for Computing Machinery, New York, NY, USA, 3561–3562. https://doi.org/10.1145/3394486.3406477
@@
- Jouppi et al. (2017) Norman P. Jouppi, Cliff Young, Nishant Patil, David Patterson, Gaurav Agrawal, Raminder Bajwa, Sarah Bates, Suresh Bhatia, Nan Boden, Al Borchers, Rick Boyle, Pierre luc Cantin, Clifford Chao, Chris Clark, Jeremy Coriell, Mike Daley, Matt Dau, Jeffrey Dean, Ben Gelb, Tara Vazir Ghaemmaghami, Rajendra Gottipati, William Gulland, Robert Hagmann, C. Richard Ho, Doug Hogberg, John Hu, Robert Hundt, Dan Hurt, Julian Ibarz, Aaron Jaffey, Alek Jaworski, Alexander Kaplan, Harshit Khaitan, Andy Koch, Naveen Kumar, Steve Lacy, James Laudon, James Law, Diemthu Le, Chris Leary, Zhuyuan Liu, Kyle Lucke, Alan Lundin, Gordon MacKean, Adriana Maggiore, Maire Mahony, Kieran Miller, Rahul Nagarajan, Ravi Narayanaswami, Ray Ni, Kathy Nix, Thomas Norrie, Mark Omernick, Narayana Penukonda, Andy Phelps, Jonathan Ross, Matt Ross, Amir Salek, Emad Samadiani, Chris Severn, Gregory Sizikov, Matthew Snelham, Jed Souter, Dan Steinberg, Andy Swing, Mercedes Tan, Gregory Thorson, Bo Tian, Horia Toma, Erick Tuttle, Vijay Vasudevan, Richard Walter, Walter Wang, Eric Wilcox, and Doe Hyun Yoon. 2017. In-Datacenter Performance Analysis of a Tensor Processing Unit. arXiv:1704.04760 [cs.AR]
@@
- Leclerc et al. (2023) Guillaume Leclerc, Andrew Ilyas, Logan Engstrom, Sung Min Park, Hadi Salman, and Aleksander Mądry. 2023. FFCV: Accelerating Training by Removing Data Bottlenecks. Proceedings of the IEEE Computer Society Conference on Computer Vision and Pattern Recognition 2023-June, 12011–12020. https://doi.org/10.1109/CVPR52729.2023.01156
- Lee et al. (2023) Claire Songhyun Lee, V Hewes, Giuseppe Cerati, Jim Kowalkowski, Adam Aurisano, Ankit Agrawal, Alok Choudhary, and Wei-Keng Liao. 2023. A Case Study of Data Management Challenges Presented in Large-Scale Machine Learning Workflows. In 2023 IEEE/ACM 23rd International Symposium on Cluster, Cloud and Internet Computing (CCGrid). 71–81. https://doi.org/10.1109/CCGrid57682.2023.00017
@@
- Lee et al. (2021) Sunwoo Lee, Qiao Kang, Kewei Wang, Jan Balewski, Alex Sim, Ankit Agrawal, Alok Choudhary, Peter Nugent, Kesheng Wu, and Wei-keng Liao. 2021. Asynchronous I/O Strategy for Large-Scale Deep Learning Applications. In 2021 IEEE 28th International Conference on High Performance Computing, Data, and Analytics (HiPC). 322–331. https://doi.org/10.1109/HiPC53243.2021.00046
- Liberty et al. (2016) Edo Liberty, Kevin Lang, and Konstantin Shmakov. 2016. Stratified Sampling Meets Machine Learning. In Proceedings of The 33rd International Conference on Machine Learning (Proceedings of Machine Learning Research, Vol. 48), Maria Florina Balcan and Kilian Q. Weinberger (Eds.). PMLR, New York, New York, USA, 2320–2329. https://proceedings.mlr.press/v48/liberty16.html
- Liu et al. (2023b) Chengjian Liu, Pengcheng Shi, Yihong Li, Lixin Liang, and Lin Lin. 2023b. Efficient Data Loading for Deep Neural Network Training. In 2023 9th International Conference on Big Data Computing and Communications (BigCom). 211–218. https://doi.org/10.1109/BIGCOM61073.2023.00036
- Liu et al. (2023a) Xuanzhe Liu, Diandian Gu, Zhenpeng Chen, Jinfeng Wen, Zili Zhang, Yun Ma, Haoyu Wang, and Xin Jin. 2023a. Rise of Distributed Deep Learning Training in the Big Model Era: From a Software Engineering Perspective. ACM Trans. Softw. Eng. Methodol. 32, 6, Article 156 (sep 2023), 26 pages. https://doi.org/10.1145/3597204
- Lozhkov et al. (2024) Anton Lozhkov, Raymond Li, Loubna Ben Allal, Federico Cassano, Joel Lamy-Poirier, Nouamane Tazi, Ao Tang, Dmytro Pykhtar, Jiawei Liu, Yuxiang Wei, Tianyang Liu, Max Tian, Denis Kocetkov, Arthur Zucker, Younes Belkada, Zijian Wang, Qian Liu, Dmitry Abulkhanov, Indraneil Paul, Zhuang Li, Wen-Ding Li, Megan Risdal, Jia Li, Jian Zhu, Terry Yue Zhuo, Evgenii Zheltonozhskii, Nii Osae Osae Dade, Wenhao Yu, Lucas Krauß, Naman Jain, Yixuan Su, Xuanli He, Manan Dey, Edoardo Abati, Yekun Chai, Niklas Muennighoff, Xiangru Tang, Muhtasham Oblokulov, Christopher Akiki, Marc Marone, Chenghao Mou, Mayank Mishra, Alex Gu, Binyuan Hui, Tri Dao, Armel Zebaze, Olivier Dehaene, Nicolas Patry, Canwen Xu, Julian McAuley, Han Hu, Torsten Scholak, Sebastien Paquet, Jennifer Robinson, Carolyn Jane Anderson, Nicolas Chapados, Mostofa Patwary, Nima Tajbakhsh, Yacine Jernite, Carlos Muñoz Ferrandis, Lingming Zhang, Sean Hughes, Thomas Wolf, Arjun Guha, Leandro von Werra, and Harm de Vries. 2024. StarCoder 2 and The Stack v2: The Next Generation. arXiv:2402.19173 [cs.SE]
@@
- Marathe et al. (2013) Aniruddha Marathe, Rachel Harris, David K. Lowenthal, Bronis R. de Supinski, Barry Rountree, Martin Schulz, and Xin Yuan. 2013. A comparative study of high-performance computing on the cloud. In Proceedings of the 22nd International Symposium on High-Performance Parallel and Distributed Computing (New York, New York, USA) (HPDC ’13). Association for Computing Machinery, New York, NY, USA, 239–250. https://doi.org/10.1145/2493123.2462919
- Mathuriya et al. (2018) Amrita Mathuriya, Deborah Bard, Peter Mendygral, Lawrence Meadows, James Arnemann, Lei Shao, Siyu He, Tuomas Kärnä, Diana Moise, Simon J. Pennycook, Kristyn Maschhoff, Jason Sewall, Nalini Kumar, Shirley Ho, Michael F. Ringenburg, Prabhat Prabhat, and Victor Lee. 2018. CosmoFlow: Using Deep Learning to Learn the Universe at Scale. In SC18: International Conference for High Performance Computing, Networking, Storage and Analysis. 819–829. https://doi.org/10.1109/SC.2018.00068
- Mattson et al. (2020) Peter Mattson, Vijay Janapa Reddi, Christine Cheng, Cody Coleman, Greg Diamos, David Kanter, Paulius Micikevicius, David Patterson, Guenther Schmuelling, Hanlin Tang, Gu-Yeon Wei, and Carole-Jean Wu. 2020. MLPerf: An Industry Standard Benchmark Suite for Machine Learning Performance. IEEE Micro 40, 2 (2020), 8–16. https://doi.org/10.1109/MM.2020.2974843
@@
- Meng (2013) Xiangrui Meng. 2013. Scalable Simple Random Sampling and Stratified Sampling. In Proceedings of the 30th International Conference on Machine Learning (Proceedings of Machine Learning Research, Vol. 28), Sanjoy Dasgupta and David McAllester (Eds.). PMLR, Atlanta, Georgia, USA, 531–539. https://proceedings.mlr.press/v28/meng13a.html
@@
- Nakama (2009) Takéhiko Nakama. 2009. Theoretical analysis of batch and on-line training for gradient descent learning in neural networks. Neurocomputing 73, 1 (2009), 151–159. https://doi.org/10.1016/j.neucom.2009.05.017 Timely Developments in Applied Neural Computing (EANN 2007) / Some Novel Analysis and Learning Methods for Neural Networks (ISNN 2008) / Pattern Recognition in Graphical Domains.
@@
- Nguyen et al. (2022) Truong Thao Nguyen, François Trahay, Jens Domke, Aleksandr Drozd, Emil Vatai, Jianwei Liao, Mohamed Wahib, and Balazs Gerofi. 2022. Why Globally Re-shuffle? Revisiting Data Shuffling in Large Scale Deep Learning. (2022), 1085–1096. https://doi.org/10.1109/IPDPS53621.2022.00109
- Nicolae et al. (2020) Bogdan Nicolae, Jiali Li, Justin M. Wozniak, George Bosilca, Matthieu Dorier, and Franck Cappello. 2020. DeepFreeze: Towards Scalable Asynchronous Checkpointing of Deep Learning Models. In 2020 20th IEEE/ACM International Symposium on Cluster, Cloud and Internet Computing (CCGRID). 172–181. https://doi.org/10.1109/CCGrid49817.2020.00-76
- Nicolae et al. (2019) Bogdan Nicolae, Adam Moody, Elsa Gonsiorowski, Kathryn Mohror, and Franck Cappello. 2019. VeloC: Towards High Performance Adaptive Asynchronous Checkpointing at Large Scale. In 2019 IEEE International Parallel and Distributed Processing Symposium (IPDPS). 911–920. https://doi.org/10.1109/IPDPS.2019.00099
@@
- Obaid et al. (2019) Hadeel S. Obaid, Saad Ahmed Dheyab, and Sana Sabah Sabry. 2019. The Impact of Data Pre-Processing Techniques and Dimensionality Reduction on the Accuracy of Machine Learning. In 2019 9th Annual Information Technology, Electromechanical Engineering and Microelectronics Conference (IEMECON). 279–283. https://doi.org/10.1109/IEMECONX.2019.8877011
@@
- Panayotov et al. (2015a) Vassil Panayotov, Guoguo Chen, Daniel Povey, and Sanjeev Khudanpur. 2015a. LibriSpeech: an ASR corpus based on public domain audio books. http://www.openslr.org/12. Accessed: February 26, 2024.
- Panayotov et al. (2015b) Vassil Panayotov, Guoguo Chen, Daniel Povey, and Sanjeev Khudanpur. 2015b. Librispeech: An ASR corpus based on public domain audio books. In 2015 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). 5206–5210. https://doi.org/10.1109/ICASSP.2015.7178964
- Paul et al. (2020) Arnab K. Paul, Olaf Faaland, Adam Moody, Elsa Gonsiorowski, Kathryn Mohror, and Ali R. Butt. 2020. Understanding HPC Application I/O Behavior Using System Level Statistics. In 2020 IEEE 27th International Conference on High Performance Computing, Data, and Analytics (HiPC). IEEE, Pune, India, 202–211. https://doi.org/10.1109/HiPC50609.2020.00034
- Paul et al. (2021) Arnab K. Paul, Ahmad Maroof Karimi, and Feiyi Wang. 2021. Characterizing Machine Learning I/O Workloads on Leadership Scale HPC Systems. In 2021 29th International Symposium on Modeling, Analysis, and Simulation of Computer and Telecommunication Systems (MASCOTS). 1–8. https://doi.org/10.1109/MASCOTS53633.2021.9614303
- Penedo et al. (2023) Guilherme Penedo, Quentin Malartic, Daniel Hesslow, Ruxandra Cojocaru, Alessandro Cappelli, Hamza Alobeidli, Baptiste Pannier, Ebtesam Almazrouei, and Julien Launay. 2023. The RefinedWeb Dataset for Falcon LLM: Outperforming Curated Corpora with Web Data, and Web Data Only. arXiv:2306.01116 [cs.CL]
@@
- Párraga et al. (2021) Edixon Párraga, Betzabeth León, Román Bond, Diego Encinas, Aprigio Bezerra, Sandra Mendez, Dolores Rexachs, and Emilio Luque. 2021. Analyzing the I/O Patterns of Deep Learning Applications. Vol. 1444 CCIS. Springer Science and Business Media Deutschland GmbH, N/A, 3–16. https://doi.org/10.1007/978-3-030-84825-5_1
@@
- Rasley et al. (2020) Jeff Rasley, Samyam Rajbhandari, Olatunji Ruwase, and Yuxiong He. 2020. DeepSpeed: System Optimizations Enable Training Deep Learning Models with Over 100 Billion Parameters. In Proceedings of the 26th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining (Virtual Event, CA, USA) (KDD ’20). Association for Computing Machinery, New York, NY, USA, 3505–3506. https://doi.org/10.1145/3394486.3406703
- Reddi et al. (2020) Vijay Janapa Reddi, Christine Cheng, David Kanter, Peter Mattson, Guenther Schmuelling, Carole-Jean Wu, Brian Anderson, Maximilien Breughe, Mark Charlebois, William Chou, Ramesh Chukka, Cody Coleman, Sam Davis, Pan Deng, Greg Diamos, Jared Duke, Dave Fick, J. Scott Gardner, Itay Hubara, Sachin Idgunji, Thomas B. Jablin, Jeff Jiao, Tom St. John, Pankaj Kanwar, David Lee, Jeffery Liao, Anton Lokhmotov, Francisco Massa, Peng Meng, Paulius Micikevicius, Colin Osborne, Gennady Pekhimenko, Arun Tejusve Raghunath Rajan, Dilip Sequeira, Ashish Sirasao, Fei Sun, Hanlin Tang, Michael Thomson, Frank Wei, Ephrem Wu, Lingjie Xu, Koichi Yamada, Bing Yu, George Yuan, Aaron Zhong, Peizhao Zhang, and Yuchen Zhou. 2020. MLPerf Inference Benchmark. arXiv:1911.02549 [cs.LG]

### v2

@@ Acknowledgements.
- et al (2022a) Assaf Eisenman et al. 2022a. Check-N-Run: a Checkpointing System for Training Deep Learning Recommendation Models. In 19th USENIX Symposium on (NSDI) 22. USENIX Association, Renton, WA, 929–943. https://www.usenix.org/conference/nsdi22/presentation/eisenman
- et al (2024) Anton Lozhkov et al. 2024. StarCoder 2 and The Stack v2: The Next Generation. arXiv:2402.19173 [cs.SE]
- et al (2021a) Alonso-Monsalve Saúl et al. 2021a. Analyzing the distributed training of deep-learning models via data locality. In 2021 29th Euromicro International Conference on Parallel, Distributed and Network-Based Processing (PDP). 117–121. https://doi.org/10.1109/PDP52278.2021.00026
- et al (2014) Ali Peshawa Jamal Muhammad et al. 2014. Data normalization and standardization: a technical report. Mach Learn Tech Rep 1, 1 (2014), 1–6.
- et al (2021b) Bez Jean Luca et al. 2021b. I/O Bottleneck Detection and Tuning: Connecting the Dots using Interactive Log Analysis. In 2021 IEEE/ACM Sixth International Parallel Data Systems Workshop (PDSW). 15–22. https://doi.org/10.1109/PDSW54622.2021.00008
- et al (2018a) Chen Chia-Yu et al. 2018a. AdaComp : Adaptive Residual Gradient Compression for Data-Parallel Distributed Training. Proceedings of the AAAI Conference on Artificial Intelligence 32, 1 (Apr. 2018). https://doi.org/10.1609/aaai.v32i1.11728
- et al (2019a) Chowdhury Fahim et al. 2019a. I/O Characterization and Performance Evaluation of BeeGFS for Deep Learning. In Proceedings of the 48th International Conference on Parallel Processing (Kyoto, Japan). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3337821.3337902
- et al (2020a) Chien Steven W. D. et al. 2020a. tf-Darshan: Understanding Fine-grained I/O Performance in Machine Learning Workloads. In 2020 IEEE International CLUSTER. IEEE Computer Society, Los Alamitos, CA, USA, 359–370. https://doi.org/10.1109/CLUSTER49012.2020.00046
- et al (2012) Chen Tianshi et al. 2012. BenchNN: On the broad potential application scope of hardware neural network accelerators. Proceedings - 2012 IEEE International Symposium on Workload Characterization, IISWC 2012, 36–45. https://doi.org/10.1109/IISWC.2012.6402898
- et al (2023b) Chen Weijian et al. 2023b. iCache: An Importance-Sampling-Informed Cache for Accelerating I/O-Bound DNN Model Training. In 2023 IEEE International Symposium on High-Performance Computer Architecture (HPCA). 220–232. https://doi.org/10.1109/HPCA56546.2023.10070964
- et al (2021c) Devarajan Hariharan et al. 2021c. DLIO: A Data-Centric Benchmark for Scientific Deep Learning Applications. In 2021 IEEE/ACM 21st International Symposium on Cluster, Cloud and Internet Computing (CCGrid). 81–91. https://doi.org/10.1109/CCGrid51090.2021.00018
- et al (2021d) Dryden Nikoli et al. 2021d. Clairvoyant Prefetching for Distributed Machine Learning I/O. In Proceedings of the International Conference for High Performance Computing, Networking, Storage and Analysis. Association for Computing Machinery. https://doi.org/10.1145/3458817.3476181
- et al (2019b) Daniel Smilkov et al. 2019b. TensorFlow.js: Machine Learning for the Web and Beyond. arXiv:1901.05350 [cs.LG]
- et al (2020b) Dey Tonmoy et al. 2020b. Optimizing Asynchronous Multi-Level Checkpoint/Restart Configurations with Machine Learning. In 2020 IEEE International Parallel and Distributed Processing Symposium Workshops (IPDPSW). 1036–1043. https://doi.org/10.1109/IPDPSW50202.2020.00174
- et al (2021e) Edixon Párraga et al. 2021e. Analyzing the I/O Patterns of Deep Learning Applications. Vol. 1444 CCIS. Springer Science and Business Media Deutschland GmbH, N/A, 3–16. https://doi.org/10.1007/978-3-030-84825-5_1
- et al (2023c) Guillaume Leclerc et al. 2023c. FFCV: Accelerating Training by Removing Data Bottlenecks. Proceedings of the IEEE Computer Society Conference on Computer Vision and Pattern Recognition 2023-June, 12011–12020. https://doi.org/10.1109/CVPR52729.2023.01156
- et al (2023d) Guilherme Penedo et al. 2023d. The RefinedWeb Dataset for Falcon LLM: Outperforming Curated Corpora with Web Data, and Web Data Only. arXiv:2306.01116 [cs.CL]
- et al (2022b) Hu Yang et al. 2022b. PipeEdge: Pipeline Parallelism for Large-Scale Model Inference on Heterogeneous Edge Devices. In 2022 25th Euromicro Conference on Digital System Design (DSD). 298–307. https://doi.org/10.1109/DSD57027.2022.00048
- et al (2020c) Jain Abhinav et al. 2020c. Overview and Importance of Data Quality for Machine Learning Tasks. In Proceedings of the 26th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining. Association for Computing Machinery. https://doi.org/10.1145/3394486.3406477
- et al (2023e) Jong Youl Choi et al. 2023e. DDStore: Distributed Data Store for Scalable Training of Graph Neural Networks on Large Atomistic Modeling Datasets. ACM International Conference Proceeding Series, 941–950. https://doi.org/10.1145/3624062.3624171
- et al (2016a) Liberty Edo et al. 2016a. Stratified Sampling Meets Machine Learning. In Proceedings of The 33rd International Conference on Machine Learning (Proceedings of Machine Learning Research, Vol. 48). PMLR, New York, New York, USA, 2320–2329. https://proceedings.mlr.press/v48/liberty16.html
- et al (2021f) Lee Sunwoo et al. 2021f. Asynchronous I/O Strategy for Large-Scale Deep Learning Applications. In 2021 IEEE 28th International Conference on High Performance Computing, Data, and Analytics (HiPC). 322–331. https://doi.org/10.1109/HiPC53243.2021.00046
- et al (2013) Marathe Aniruddha et al. 2013. A comparative study of high-performance computing on the cloud. In Proceedings of the 22nd International Symposium on HPDC. Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/2493123.2462919
- et al (2018b) Mathuriya Amrita et al. 2018b. CosmoFlow: Using Deep Learning to Learn the Universe at Scale. In SC18: International Conference for High Performance Computing, Networking, Storage and Analysis. 819–829. https://doi.org/10.1109/SC.2018.00068
- et al (2020d) Mattson Peter et al. 2020d. MLPerf: An Industry Standard Benchmark Suite for Machine Learning Performance. IEEE Micro 40, 2 (2020), 8–16. https://doi.org/10.1109/MM.2020.2974843
- et al (2019c) Nicolae Bogdan et al. 2019c. VeloC: Towards High Performance Adaptive Asynchronous Checkpointing at Large Scale. In 2019 IEEE International Parallel and Distributed Processing Symposium (IPDPS). 911–920. https://doi.org/10.1109/IPDPS.2019.00099
- et al (2020e) Nicolae Bogdan et al. 2020e. DeepFreeze: Towards Scalable Asynchronous Checkpointing of Deep Learning Models. In 2020 20th IEEE/ACM International Symposium on Cluster, Cloud and Internet Computing (CCGRID). 172–181. https://doi.org/10.1109/CCGrid49817.2020.00-76
- et al (2017) Norman P. Jouppi et al. 2017. In-Datacenter Performance Analysis of a Tensor Processing Unit. arXiv:1704.04760 [cs.AR]
- et al (2022c) Nguyen Truong Thao et al. 2022c. Why Globally Re-shuffle? Revisiting Data Shuffling in Large Scale Deep Learning. (2022), 1085–1096. https://doi.org/10.1109/IPDPS53621.2022.00109
- et al (2020f) Paul Arnab K. et al. 2020f. Understanding HPC Application I/O Behavior Using System Level Statistics. In 2020 IEEE 27th International Conference on High Performance Computing, Data, and Analytics (HiPC). IEEE, Pune, India, 202–211. https://doi.org/10.1109/HiPC50609.2020.00034
- et al (2021g) Paul Arnab K et al. 2021g. Characterizing Machine Learning I/O Workloads on Leadership Scale HPC Systems. In 2021 29th International Symposium MASCOTS. 1–8. https://doi.org/10.1109/MASCOTS53633.2021.9614303
- et al (2022d) Pablo Villalobos et al. 2022d. Machine Learning Model Sizes and the Parameter Gap. arXiv:2207.02852 [cs.LG]
- et al (2016b) Robert Adolf et al. 2016b. Fathom: reference workloads for modern deep learning methods. 2016 IEEE International Symposium on Workload Characterization (IISWC) (2016), 1–10. https://api.semanticscholar.org/CorpusID:1809844
- et al (2020g) Rasley Jeff et al. 2020g. In Proceedings of the 26th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining (Virtual Event, CA, USA) (KDD ’20). Association for Computing Machinery, New York, NY, USA, 3505–3506. https://doi.org/10.1145/3394486.3406703
- et al (2021h) Steven Farrell et al. 2021h. MLPerf HPC: A Holistic Benchmark Suite for Scientific Machine Learning on HPC Systems. arXiv:2110.11466 [cs.LG] https://arxiv.org/abs/2110.11466
- et al (2016c) Snyder Shane et al. 2016c. Modular HPC I/O Characterization with Darshan. In 2016 5th Workshop on Extreme-Scale Programming Tools (ESPT). 2016 5th workshop on extreme-scale programming tools (ESPT), Salt Lake City, UT, USA, 9–17. https://doi.org/10.1109/ESPT.2016.006
- et al (2020h) Smith Samuel et al. 2020h. On the Generalization Benefit of Noise in Stochastic Gradient Descent. In Proceedings of the 37th International Conference on Machine Learning, Hal Daumé III and Aarti Singh (Eds.), Vol. 119. PMLR, Cornell University. https://proceedings.mlr.press/v119/smith20a.html
- et al (2022e) Sun Xiaoyang et al. 2022e. STRONGHOLD: Fast and Affordable Billion-Scale Deep Learning Model Training. In SC22: International Conference for High Performance Computing, Networking, Storage and Analysis. 1–17. https://doi.org/10.1109/SC41404.2022.00076
- et al (2020i) Tom B. Brown et al. 2020i. Language Models are Few-Shot Learners. arXiv:2005.14165 [cs.CL]
- et al (2020j) Vijay Janapa Reddi et al. 2020j. MLPerf Inference Benchmark. arXiv:1911.02549 [cs.LG]
- et al (2021i) Weihua Hu et al. 2021i. Open Graph Benchmark: Datasets for Machine Learning on Graphs Steering Committee. https://ogb.stanford.edu.
- et al (2020k) Wan Lipeng et al. 2020k. I/O Performance Characterization and Prediction through Machine Learning on HPC Systems. In CUG2020. CUG, Dallas Convention Center Room A309-A310 650 S Griffin St Dallas, TX 75202, 10.
- et al (2022f) Xie Jian et al. 2022f. A Deep Learning Dataloader with Shared Data Preparation. In Advances in Neural Information Processing Systems, S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Eds.), Vol. 35. Curran Associates, Inc., Virtual, 17146–17156.
- et al. (2020) Yao S. et al. 2020. Deep compressive offloading: speeding up neural network inference by trading edge computation for network latency. In Proc. 18th SenSys (SenSys ’20). Association for Computing Machinery, New York, NY, USA, 476–488. https://doi.org/10.1145/3384419.3430898
- et al (2018c) Zhu Maohua et al. 2018c. Performance Evaluation and Optimization of HBM-Enabled GPU for Data-Intensive Applications. IEEE Transactions on Very Large Scale Integration (VLSI) Systems 26, 5 (2018), 831–840. https://doi.org/10.1109/TVLSI.2018.2791442
- et al (2018d) Zhu Yue et al. 2018d. Entropy-Aware I/O Pipelining for Large-Scale Deep Learning on HPC Systems. In 2018 IEEE 26th International Symposium on Modeling, Analysis, and Simulation of Computer and Telecommunication Systems (MASCOTS). 145–156. https://doi.org/10.1109/MASCOTS.2018.00023
@@
- Google Cloud (2024a) Google Cloud. 2024a. Google Analytics Data on BigQuery. https://cloud.google.com/bigquery/public-data/google-analytics Accessed: 2024-09-12.
- Google Cloud (2024b) Google Cloud. 2024b. Google BigQuery Public Datasets. https://cloud.google.com/bigquery/public-data Accessed: 2024-09-12.
@@
- Jie et al. (2023) Fei Jie, Haijun Zhang, Jinxiang Wang, and Zhenhua Yu. 2023. ASRDataset: A Multi-granularity Shuffle System for Preparing Large-scale ASR Training Data. In 2023 IEEE International Conference on Knowledge Graph (ICKG). 69–76. https://doi.org/10.1109/ICKG59574.2023.00014
- Jin et al. (2024) Hai Jin, Zhanyang Zhu, Ligang He, Yuhao Li, Yusheng Hua, and Xuanhua Shi. 2024. MMDataLoader: Reusing Preprocessed Data Among Concurrent Model Training Tasks. IEEE Trans. Comput. 73, 2 (2024), 510–522. https://doi.org/10.1109/TC.2023.3336161
@@
- Lee et al. (2008) Choonghwan Lee, MuQun Yang, and Ruth Aydt. 2008. NetCDF-4 Performance Report. (2008), 22 pages. https://www.hdfgroup.org/pubs/papers/2008-06_netcdf4_perf_report.pdf
- Lee (2023) Claire Songhyun et al Lee. 2023. A Case Study of Data Management Challenges Presented in Large-Scale Machine Learning Workflows. In 2023 IEEE/ACM 23rd International Symposium on Cluster, Cloud and Internet Computing (CCGrid). 71–81. https://doi.org/10.1109/CCGrid57682.2023.00017
@@
- Li et al. (2020) En Li, Liekang Zeng, Zhi Zhou, and Xu Chen. 2020. Edge AI: On-Demand Accelerating Deep Neural Network Inference via Edge Computing. IEEE Transactions on Wireless Communications 19, 1 (2020), 447–457. https://doi.org/10.1109/TWC.2019.2946140
- Liu et al. (2023) Chengjian Liu, Pengcheng Shi, Yihong Li, Lixin Liang, and Lin Lin. 2023. Efficient Data Loading for Deep Neural Network Training. In 2023 9th International Conference on Big Data Computing and Communications (BigCom). 211–218. https://doi.org/10.1109/BIGCOM61073.2023.00036
- Liu et al. (2020) Xiaolan Liu, Jiadong Yu, Jian Wang, and Yue Gao. 2020. Resource Allocation With Edge Computing in IoT Networks via Machine Learning. IEEE Internet of Things Journal 7, 4 (2020), 3415–3426. https://doi.org/10.1109/JIOT.2020.2970110
@@
- Mammen (2021) Priyanka Mary Mammen. 2021. Federated Learning: Opportunities and Challenges. arXiv:2101.05428 [cs.LG] https://arxiv.org/abs/2101.05428
- Maurya et al. (2024) Avinash Maurya et al. 2024. DataStates-LLM: Lazy Asynchronous Checkpointing for Large Language Models. In Proceedings of the 33rd International Symposium on High-Performance Parallel and Distributed Computing (HPDC ’24). ACM, 227–239. https://doi.org/10.1145/3625549.3658685
@@
- McMahan et al. (2023) H. Brendan McMahan, Eider Moore, Daniel Ramage, Seth Hampson, and Blaise Agüera y Arcas. 2023. Communication-Efficient Learning of Deep Networks from Decentralized Data. arXiv:1602.05629 [cs.LG] https://arxiv.org/abs/1602.05629
@@
- Meng (2013) Xiangrui Meng. 2013. Scalable Simple Random Sampling and Stratified Sampling. In Proceedings of the 30th International Conference on Machine Learning, Sanjoy Dasgupta and David McAllester (Eds.), Vol. 28. PMLR, Atlanta, Georgia, USA. https://proceedings.mlr.press/v28/meng13a.html
@@
- Microsoft Azure (2024a) Microsoft Azure. 2024a. Azure Open Datasets. https://azure.microsoft.com/en-us/services/open-datasets/ Accessed: 2024-09-12.
- Microsoft Azure (2024b) Microsoft Azure. 2024b. NYC Taxi & Limousine Commission Data. https://azure.microsoft.com/en-us/services/open-datasets/nyc-taxi-limousine/ Accessed: 2024-09-12.
- Murray et al. (2021) Derek G. Murray, Jiri Simsa, Ana Klimovic, and Ihor Indyk. 2021. tf.data: A Machine Learning Data Processing Framework. arXiv:2101.12127 [cs.LG]
- Nakama (2009) Takéhiko Nakama. 2009. Theoretical analysis of batch and on-line training for gradient descent learning in neural networks. Neurocomputing 73, 1 (2009), 151–159. https://doi.org/10.1016/j.neucom.2009.05.017
@@
- Obaid et al. (2019) Hadeel S. Obaid, Saad Ahmed Dheyab, and Sana Sabah Sabry. 2019. The Impact of Data Pre-Processing Techniques and Dimensionality Reduction on the Accuracy of Machine Learning. In 2019 9th Annual IEMECON. 279–283. https://doi.org/10.1109/IEMECONX.2019.8877011
@@
- Panayotov et al. (2015) Vassil Panayotov, Guoguo Chen, Daniel Povey, and Sanjeev Khudanpur. 2015. LibriSpeech: an ASR corpus based on public domain audio books. http://www.openslr.org/12. Accessed: February 26, 2024.
- Petersen (2015) Torben Kling Petersen. 2015. Inside the Lustre File System. Technical Paper. SEAGATE Technology.
@@
- Schimmelpfennig et al. (2021) Frederic Schimmelpfennig et al. 2021. Streamlining distributed Deep Learning I/O with ad hoc file systems. In 2021 IEEE International Conference on Cluster Computing (CLUSTER). 169–180. https://doi.org/10.1109/Cluster48925.2021.00062
@@
- Serizawa et al. (2019) Kazuhiro Serizawa et al. 2019. Accelerating Machine Learning I/O by Overlapping Data Staging and Mini-batch Generations. In Proc. 6th IEEE/ACM BDCAT (Auckland, New Zealand) (BDCAT ’19). Association for Computing Machinery, 31–34. https://doi.org/10.1145/3365109.3368768